Brad wants to buy flowers for his friend with $45. The daisies are $1 each and the roses are $2 each. He buys 3 more daisies than roses. How much did the roses cost?
Answer by Maths68(1474)

Let
Brad buys Roses = x
Brad buys Daisies = x+3
Given
1 Rose cost = $2
1 Daisy = $1
Total cost of flowers = 45
Roses*cost of a Rose + Daisies * cost of a Daisy = Total Cost of flowers
x*2+(x+3)*1=45
2x+x+3=45
3x+3=45
3x=45-3
3x=42
3x/3=42/3
x=14
Brad buys Roses = x = 14
Cost of Roses = 14*2=$28

Check
======
Cost of 14 roses = 28
Cost of 17 Daisies = 17
--------------------------
Cost of Flowers = $45